OpenStudy (anonymous):

check my answer? What is the slope between the two points (6, 20) and (2, 4)? 1/4 -1/4 4 -4 i got 4

OpenStudy (uri):

Slope formula :(y2-y1/x2-x1) =4-20/2-6 =-16/-4 =-4 Ans.

OpenStudy (uri):

Wait 4 ans - will be cancelled.

OpenStudy (anonymous):

okay thank you (: @uri

OpenStudy (anonymous):

\[\frac{4-20}{2-6}=4 \]